# Electrically Tunable and Enhanced Nonlinearity of Moiré Exciton-Polaritons in Transition Metal Dichalcogenide Bilayers

### v2

@@ Abstract
In this Letter we develop a microscopic theory of twisting-dependent optical response of 2D bilayers, taking into account a rich structure of moiré modes. Solving a Wannier equation within the multi-Gaussian basis expansion, we resolve excitonic modes emerging from hybridization of interlayer and intralayer states at varying twisting angles and external bias. Crucially, this allows calculating nonlinear scattering for different modes without simplifying their structure, as well as describing strong light-matter coupling. We observe that the Umklapp scattering involving neighboring mini Brillouin zones (mBZ) of the moiré lattice play an important role for nonlinearity, leading to its enhancement for non-zero $\theta$. For bilayers with small twisting we reveal an attractive nonlinearity driven by exchange processes, and recover a conventional dipolar repulsion for large ($>3^{\circ}$) twists. Our results can offer microscopic insights into nonlinear redshifts measured experimentally [30], and open avenues for quantitative studies of quantum moiré polaritons.
Model.—We consider a bilayer Hamiltonian $\mathcal{H}=\mathcal{H}_{0}+\mathcal{H}_{\mathrm{int}}+\mathcal{H}_{t}$, consisting of the free energy terms, Coulomb interaction term, and the tunneling term, respectively. The free energy for electrons (e) and hole (h) in different layers reads as $\mathcal{H}_{0}=\sum_{\mathbf{k}\sigma}\sum_{\ell=1,2}[\varepsilon^{c}_{\ell\sigma}(\mathbf{k})a^{\ell,\dagger}_{\mathbf{k}\sigma}a^{\ell}_{\mathbf{k}\sigma}+\varepsilon^{v}_{\ell\sigma}(\mathbf{k})b^{\ell,\dagger}_{\mathbf{k}\sigma}b^{\ell}_{\mathbf{k}\sigma}]$, where a crystal momentum $\mathbf{k}$ is measured from the $\Gamma$-point in each layer labeled as $\ell=1,2$, and $\sigma=\uparrow,\downarrow$ is the spin index. The density-density interaction term reads $\mathcal{H}_{\mathrm{int}}=\frac{1}{2L^{2}}\sum_{\ell\ell^{\prime},\mathbf{q}}v_{\ell\ell^{\prime}}(\mathbf{q})\rho_{\ell\mathbf{q}}\rho_{\ell^{\prime},-\mathbf{q}}$, where $L^{2}$ is a sample area. Since only the low-energy e-h pairs are relevant for forming the exciton-bound state [47, 48], we expand the dispersion near the valleys at $\mathbf{K}_{\ell}$ (band edge) as $\varepsilon^{\mathrm{b}}_{\ell\sigma}(\mathbf{k})=\Delta^{\mathrm{b}}_{\ell\sigma}\pm(\mathbf{k}-\mathbf{K}_{\ell})^{2}/(2m_{\ell}^{\mathrm{b}})$ with $\mathrm{b}=c,v$ being the conduction and valance band index, and $m^{\mathrm{b}}_{\ell}$ being the $\mathrm{b}$-band mass in layer $\ell$. We let the energy offset for the $\mathrm{b}$-band edge as $\Delta^{\mathrm{b}}_{\ell\sigma}$ that is measured from the topmost valance band (see Fig. 1c). In the Coulomb term, $\rho_{\ell\mathbf{q}}=\sum_{\mathbf{k}\sigma}(a^{\ell\dagger}_{\mathbf{k}+\mathbf{q},\sigma}a^{\ell}_{\mathbf{k}\sigma}+b^{\ell\dagger}_{\mathbf{k}+\mathbf{q},\sigma}b^{\ell}_{\mathbf{k}\sigma})$ is the charge density, operator and the screened potential between electrons is of the Keldysh-Rytova form [49],
@@
with $\Delta\mathbf{K}=\mathbf{K}_{\ell_{c}}-\mathbf{K}_{\ell_{v}}$, $\varepsilon^{s}_{\sigma}=\Delta^{c}_{\ell_{c}\sigma}-\Delta^{v}_{\ell_{v}\sigma}$, and $E^{s}_{\alpha\sigma}(\mathbf{Q})$ being the exciton energy. The reduced mass and total mass are $\mu_{s}=(1/m^{c}_{\ell_{c}}+1/m^{v}_{\ell_{v}})^{-1}$ and $M_{s}=m^{c}_{\ell_{c}}+m^{v}_{\ell_{v}}$, respectively. The state (3) and associated wavefunction will serve as the basis for studying the hybridized exciton.
@@
with $|0\rangle$ being the ground state of the bilayer. In Eq. (7), we label the hybridized moiré exciton eigenstates by index $\bar{\alpha}$ with a bar. More discussion of Eq. (7) can be found in Supplemental Material (SM) [52] and Ref. [50].
@@
where the direct [36] and exchange interactions are
@@
that are depicted in Fig. 1d. Here, $\mathbf{q}$ is the transferred momentum between $\bar{\alpha}$- and $\bar{\beta}$-exciton and the scattering potential is $\Gamma_{\nu\tilde{\nu}}^{\nu^{\prime}\tilde{\nu}^{\prime}}(\mathbf{k}\tilde{\mathbf{k}},\mathbf{q})=\sum_{\ell\ell^{\prime}}f^{\ell}_{\nu\sigma}(\mathbf{k},\mathbf{q})(v_{\ell\ell^{\prime}}(\mathbf{q})/2L^{2})f^{\ell^{\prime}}_{\tilde{\nu}\sigma}(\tilde{\mathbf{k}},-\mathbf{q})\phi^{\ast}_{\nu^{\prime}}(\mathbf{k})\phi^{\ast}_{\tilde{\nu}^{\prime}}(\tilde{\mathbf{k}})$, with excitonic wavefunction being expressed in $\nu$-index notation as $\phi_{\nu\sigma}(\mathbf{k})=\phi^{s}_{\alpha\sigma}(\mathbf{k}-m^{c}_{\ell_{c}}/M_{s}\bm{g})$, and the factor $f^{\ell}_{\nu\sigma}(\mathbf{k},\mathbf{q})=\delta_{\ell_{c}\ell}\phi_{\nu\sigma}(\mathbf{k}-\mathbf{q})-\delta_{\ell_{v}\ell}\phi_{\nu\sigma}(\mathbf{k})$. The calculation of the direct interaction $U_{\nu\tilde{\nu}}^{\nu^{\prime}\tilde{\nu}^{\prime}}$ is straightforward while evaluating the exchange interaction $V_{\nu\tilde{\nu}}^{\nu^{\prime}\tilde{\nu}^{\prime}}$ is rather involved and is detailed in SM [52].
@@
Figure 2: (a) Exciton energy $\mathcal{E}$ ($\sigma=\uparrow$) and twisting angle $\theta$ dependence. The points are color coded by its light-matter coupling strength $g_{c}$. The hybridized excitons weakly interact with light are shown in gray dots. (Inset) the moiré exciton spectrum is calculated by considering the Umklapp scattering that hybridizes the exciton in zone 1 (purple area) and zone 2 (gray area). In the weak Umklapp region, the interlayer and intralayer exciton hybridized with each mBZ separately. (b) $\theta$-dependence of the exciton-exciton interaction. (Inset) the interaction (wavy black curves) between interlayer excitons with electron exchange (red dashed curves).
@@
$\mathcal{H}_{\mathrm{xp}}=$ $\omega_{c}c^{\dagger}_{0}c_{0}+\sum_{\bar{\alpha}}(\mathcal{E}^{\bar{\alpha}}_{\sigma}+\tfrac{1}{L^{2}}g^{\bar{\alpha}}_{X}\hat{x}^{\dagger}_{\bar{\alpha}\sigma}\hat{x}_{\bar{\alpha}\sigma})\hat{x}^{\dagger}_{\bar{\alpha}\sigma}\hat{x}_{\bar{\alpha}\sigma}$
$+\sum_{\bar{\alpha}}g^{\bar{\alpha}}_{c}(c^{\dagger}_{0}\hat{x}_{\bar{\alpha}\sigma}+\mathrm{h.c.})$ (10)
where the X-X interaction is $g^{\bar{\alpha}}_{X}\approx L^{2}\Delta^{\bar{\alpha}\bar{\alpha}}$. The exciton-photon interaction is $g_{c}^{\bar{\alpha}}=\frac{1}{2}\sum_{\alpha\ell}\Omega^{\ell}_{\alpha}\bar{C}^{\bar{\alpha},\ell\ell}_{\alpha\sigma}$, where $\Omega^{\ell}_{\alpha}=2\sum_{\mathbf{k}}d^{\ell}_{cv}\phi^{\ell\ell}_{\alpha}(\mathbf{k})$ is the Rabi splitting of the exciton in $\ell$-th layer and state $\alpha$. Therefore, we constructed the microscopic model of twisting-dependent moiré polaritons with tunable nonlinearity $g^{\bar{\alpha}}_{X}$ and light-matter coupling $g_{c}$.
@@
Result and discussion.—To demonstrate the tunability of moiré polaritons described by Eq. (Electrically Tunable and Enhanced Nonlinearity of Moiré Exciton-Polaritons in Transition Metal Dichalcogenide Bilayers), we investigate the band hybridization effects for a MoS2 bilayer marginally twisted from anti-parallel stacking/H-type ($\theta\approx 60^{\circ}$) encapsulated by hexagonal boron nitride ($\epsilon=4$). We adopt the band parameters from Ref. [60] and use interlayer hoppings of $t_{c}(\mathbf{K}_{\ell})=2.1$ meV and $t_{v}(\mathbf{K}_{\ell})=14.5$ meV [51]. Using the Gaussian basis function expansion [48, 61, 62], we solve Eqs. (Electrically Tunable and Enhanced Nonlinearity of Moiré Exciton-Polaritons in Transition Metal Dichalcogenide Bilayers) and (7), and plot the $\theta$-dependence of exciton energies $\mathcal{E}^{\bar{\alpha}}_{\sigma}$ in Fig. 2a. The corresponding strength of nonlinearity $g^{\bar{\alpha}}_{X}$ is shown in Fig. 2b. In the plots each point corresponds to moiré excitons with an associated index $\bar{\alpha}$, and their light-matter coupling strength ($g_{c}^{\bar{\alpha}}$) is color-coded. In Fig. 2a, we can see two lower branches of bright excitonic modes that correspond to the hybridized intralayer exciton hX ($\mathcal{E}\approx 1.98$ eV) and hybridized interlayer exciton hIX ($\mathcal{E}\approx 2.05$ eV). These results match previous calculations for single-particle properties discussed in Ref. [50].
@@
In our analysis we focus on the nonlinear response and X-X interaction strength shown in Fig. 2b. In the plot we keep only the dominant contributions from Eq. (9) with $\bm{g}^{\prime}=\bm{g}$ for simplicity, since the Coulomb scattering processes with nonzero momentum transfer between the excitons are rapidly suppressed as $\theta\gtrsim 0.5^{\circ}$. First of all, we find an enhancement in the nonlinear interaction at the small-$\theta$ regime indicating the strong influence of the band hybridization due to the Umklapp processes. We find strong attractive nonlinearity for hIX while the nonlinear interaction for hX remains repulsive. Intriguingly, these results correlate with measurements in MoS2 bilayers, where redshifts were observed [30]. The emergence of attractive nonlinearity may be understood as a manifestation of strong anisotropic nature of 2D bilayers. This leads to a weaker interlayer e-h interaction as compared to the intralayer e-e and h-h interaction (Fig. 2b, inset). Hence, the total nonlinearity from each of these interacting channels does not cancel leading to stronger attractive nonlinearity. This is different from the monolayer exciton where large cancellations [61, 62, 63] between these channels result in a weaker repulsive nonlinearity ($\sim 0.5$ nm2eV). This anisotropic electronic property is unique for 2D materials that enable the realization of attractive interaction between dipolar excitons. We also note that the strong attractive X-X interaction between hIX in the small $\theta$ regime comes from the exchange scattering involving the excitonic modes in Zone 2 (Fig. 2a, inset). For instance, in the scattering processes with $\bm{g}\neq\tilde{\bm{g}}$ (Fig. 1d), the second term in the last line in Eq.(9) (all particle exchange process, repulsive) is strongly suppressed due to large momentum transfer ($\mathbf{q}=\bm{g}-\tilde{\bm{g}}$), while the exchange interaction $V_{\nu\tilde{\nu}}^{\nu^{\prime}\tilde{\nu}^{\prime}}$ and $V_{\nu\tilde{\nu}}^{\tilde{\nu}^{\prime}\nu^{\prime}}$ remain large. This makes a striking difference between moiré bilayer and untwisted bilayer ($\theta=0$ and $60^{\circ}$) where the scattering processes only take place within Zone 1 with $\bm{g}=\tilde{\bm{g}}=0$. Furthermore, in our calculation, we find weak repulsive direct interaction for hIX in a homobilayer with $\theta\approx 0$ (see SM[52]). This may be expected since a large electrical dipole moment is forbidden due to the approximate inversion symmetry. As a result, the dominant contribution to $g_{X}$ in Fig. 2b comes from the (attractive) exchange interactions. However, this can change for heterobilayers or electrically-biased samples.
In the presence of electric field $\mathbf{E}_{z}$, the dipole moment of the interlayer exciton couples to $\mathbf{E}_{z}$ giving rise to the energy Stark shift. This changes the hybridization content of hX and hIX leading to electrically-tunable optical properties. The Stark effect can be incorporated into Eq. (7) by modifying the interlayer exciton energy $E_{\nu\sigma}(\bar{\mathbf{Q}})\to E_{\nu\sigma}(\bar{\mathbf{Q}})+\mathbf{E}_{z}\cdot\bm{\mu}^{s}$, where the dipole moment $\bm{\mu}^{s}=\pm\mu_{z}\hat{\mathbf{z}}$ for interlayer exciton with $\mu_{z}\sim 5$eÅ [64, 29], and $\mu_{z}=0$ for intralayer exciton. In Figs. 3a and 3b, we calculate the absorption spectrum using the Fermi golden rule with a Lorentz broadening ($\beta=5$ meV), leading to
$\Gamma(\mathcal{E})=\frac{2\pi}{\hbar}\sum_{\bar{\alpha}}\frac{(\beta/\pi)|\langle 0|\mathcal{X}^{\bar{\alpha}}_{\sigma}\mathcal{H}_{\mathrm{sc}}c^{\dagger}_{\mathbf{Q}}|0\rangle|^{2}}{[\mathcal{E}^{\bar{\alpha}}_{\sigma}+n_{\bar{\alpha}}g^{\bar{\alpha}}_{X}-\mathcal{E}(\mathbf{Q})]^{2}+\beta^{2}},$ (11)
@@
where the photon dispersion is $\mathcal{E}(\mathbf{Q})=\hbar|\mathbf{Q}|$. In the absorption, we also take into account the nonlinear energy shift due to the background excitons with density $n_{\bar{\alpha}}=\langle\hat{x}^{\dagger}_{\bar{\alpha}\sigma}\hat{x}_{\bar{\alpha}\sigma}\rangle/L^{2}$ (pump-power dependent). Two brightest branches independent of $E_{z}$ correspond to the hybridized A and B intralayer excitons. The other dimmer branches with strong response to $E_{z}$ are the hIXs. Following one hIX branch indicated by the red and blue dots in Figs. 3a and 3b, we can see that the absorption peaks are enhanced as the hIX approaches the hX branches since hIX gains more intralayer exciton component. Furthermore, we also observe that the hIX nonlinear redshift turns into a blueshift as $E_{z}$ changes. This is partly due to the repulsive nonlinearity from the intralayer exciton component, and another contribution is coming from the polarization of hIX in the high-field regime. The later contribution enhances the repulsive dipole-dipole direct interaction of hIX. These results in Figs. 2 and 3 demonstrate that the nonlinear optical response of moiré material is electrically tunable.
@@
Next, we plot the polaritonic spectrum in Fig. 3(c) by evaluating the photonic density of states $\rho_{c}(\omega)$ of Eq. (Electrically Tunable and Enhanced Nonlinearity of Moiré Exciton-Polaritons in Transition Metal Dichalcogenide Bilayers), which is proportional to the cavity transmission [52]. Tuning the hIX energy closer to the hX energy at $E_{z}=0.5$ eV/nm, we enhance the light-matter coupling of hIX leading to the larger Rabi splitting. This also changes the hIX attractive nonlinearity (red arrows) at $E_{z}=0$ to a repulsive nonlinearity (blue arrows) with blueshift at $E_{z}=0.5$ V/nm.
@@
Conclusion and discussion.—We developed a microscopic theory for hybridized moiré excitons in twisted bilayers, and studied their nonlinear optical response in the strong light-matter coupling regime. We revealed that hybridization between layers can enhance significantly the nonlinearity of moiré excitons and polaritons, stemming from Umklapp processes for small twisting angles. Intriguingly, we find attractive nonlinear interaction for hybridized interlayer excitons (hIX), which can be tuned into repulsive by applying an external electric field. This makes moiré polariton lattices at small twisting angles an excellent platform for studying many-body effects.
@@
Note that in this study we limited ourselves to not-too-large $\theta$ regime, since the hybridization between folded bands in the mini Brillouin zone is weak. On the other hand, we do not investigate $\theta\lesssim 0.5^{\circ}$, as this is typically prevented by the lattice reconstruction [65, 66, 67, 68, 69] and also requires accounting for more Umklapp scatterings beyond those pictured in Fig. 2a. We also stress that the nonlinear properties in our study are valid for low density regimes, and studying higher-order correction is an interesting avenue for future research. This will allow describing a crossover to the strongly correlated regime close to the metal-insulator transition [70, 71, 72, 73]. We also note that while results are shown for a selected TMD configurations, our calculations for other materials and stacking confirmed qualitatively similar results. This confirms that the described moiré physics is representative of twisting processes and is rather universal.

### v3

@@ Abstract
In this Letter, we develop a microscopic theory of twisting-dependent optical response of 2D bilayers, taking into account a rich structure of moiré modes. Solving a Wannier equation within the multi-Gaussian basis expansion, we resolve excitonic modes emerging from hybridization of interlayer and intralayer states at varying twisting angles and external bias. Crucially, this allows calculating nonlinear scattering for different modes without simplifying their structure, as well as describing strong light-matter coupling. We observe that the Umklapp scattering involving neighboring mini Brillouin zones (mBZ) of the moiré lattice play an important role for nonlinearity, leading to its enhancement for non-zero $\theta$. For bilayers with small twisting we reveal an attractive nonlinearity driven by exchange processes, and recover a conventional dipolar repulsion for large ($>3^{\circ}$) twists. Our results can offer microscopic insights into nonlinear redshifts measured experimentally [30, 47], and open avenues for quantitative studies of quantum moiré polaritons.
Model.—We consider a bilayer Hamiltonian $\mathcal{H}=\mathcal{H}_{0}+\mathcal{H}_{\mathrm{int}}+\mathcal{H}_{t}$, consisting of the free energy terms, Coulomb interaction term, and the tunneling term, respectively. The free energy for electrons (e) and hole (h) in different layers reads as $\mathcal{H}_{0}=\sum_{\mathbf{k}\sigma}\sum_{\ell=1,2}[\varepsilon^{c}_{\ell\sigma}(\mathbf{k})a^{\ell,\dagger}_{\mathbf{k}\sigma}a^{\ell}_{\mathbf{k}\sigma}+\varepsilon^{v}_{\ell\sigma}(\mathbf{k})b^{\ell,\dagger}_{\mathbf{k}\sigma}b^{\ell}_{\mathbf{k}\sigma}]$, where a crystal momentum $\mathbf{k}$ is measured from the $\Gamma$-point in each layer labeled as $\ell=1,2$, and $\sigma=\uparrow,\downarrow$ is the spin index. The density-density interaction term reads $\mathcal{H}_{\mathrm{int}}=\frac{1}{2L^{2}}\sum_{\ell\ell^{\prime},\mathbf{q}}v_{\ell\ell^{\prime}}(\mathbf{q})\rho_{\ell\mathbf{q}}\rho_{\ell^{\prime},-\mathbf{q}}$, where $L^{2}$ is a sample area. Since only the low-energy e-h pairs are relevant for forming the exciton-bound state [48, 49], we expand the dispersion near the valleys at $\mathbf{K}_{\ell}$ (band edge) as $\varepsilon^{\mathrm{b}}_{\ell\sigma}(\mathbf{k})=\Delta^{\mathrm{b}}_{\ell\sigma}\pm(\mathbf{k}-\mathbf{K}_{\ell})^{2}/(2m_{\ell}^{\mathrm{b}})$ with $\mathrm{b}=c,v$ being the conduction and valence band index, and $m^{\mathrm{b}}_{\ell}$ being the $\mathrm{b}$-band mass in layer $\ell$. We let the energy offset for the $\mathrm{b}$-band edge as $\Delta^{\mathrm{b}}_{\ell\sigma}$ that is measured from the topmost valence band (see Fig. 1c). In the Coulomb term, $\rho_{\ell\mathbf{q}}=\sum_{\mathbf{k}\sigma}(a^{\ell\dagger}_{\mathbf{k}+\mathbf{q},\sigma}a^{\ell}_{\mathbf{k}\sigma}+b^{\ell\dagger}_{\mathbf{k}+\mathbf{q},\sigma}b^{\ell}_{\mathbf{k}\sigma})$ is the charge density, operator and the screened potential between electrons is of the Keldysh-Rytova form [50],
@@
with $\Delta\mathbf{K}=\mathbf{K}_{\ell_{c}}-\mathbf{K}_{\ell_{v}}$, $\varepsilon^{s}_{\sigma}=\Delta^{c}_{\ell_{c}\sigma}-\Delta^{v}_{\ell_{v}\sigma}$, and $E^{s}_{\alpha\sigma}(\mathbf{Q})$ being the exciton energy. The total mass and reduced mass are $M_{s}=m^{c}_{\ell_{c}}+m^{v}_{\ell_{v}}$, and $\mu_{s}=m^{c}_{\ell_{c}}m^{v}_{\ell_{v}}/M_{s}$.
@@
with $|0\rangle$ being the ground state. In Eq. (7), we label the hybridized moiré exciton eigenstates by index $\bar{\alpha}$ with a bar, see details in Supplemental Material (SM) [53]. The excitonic model here is different from the moiré potential approach[5, 8, 41] which does not account for the interlayer-intralayer exciton hybridization.
@@
where the direct [36] and exchange interactions[54] are
@@
that are depicted in Fig. 1d. Here, $\mathbf{q}$ is the transferred momentum between $\bar{\alpha}$- and $\bar{\beta}$-exciton and the scattering potential is $\Gamma_{\nu\tilde{\nu}}^{\nu^{\prime}\tilde{\nu}^{\prime}}(\mathbf{k}\tilde{\mathbf{k}},\mathbf{q})=\sum_{\ell\ell^{\prime}}f^{\ell}_{\nu\sigma}(\mathbf{k},\mathbf{q})(v_{\ell\ell^{\prime}}(\mathbf{q})/2L^{2})f^{\ell^{\prime}}_{\tilde{\nu}\sigma}(\tilde{\mathbf{k}},-\mathbf{q})\phi^{\ast}_{\nu^{\prime}}(\mathbf{k})\phi^{\ast}_{\tilde{\nu}^{\prime}}(\tilde{\mathbf{k}})$, with excitonic wavefunction being expressed in $\nu$-index notation as $\phi_{\nu\sigma}(\mathbf{k})=\phi^{s}_{\alpha\sigma}(\mathbf{k}-m^{c}_{\ell_{c}}/M_{s}\bm{g})$, and the factor $f^{\ell}_{\nu\sigma}(\mathbf{k},\mathbf{q})=\delta_{\ell_{c}\ell}\phi_{\nu\sigma}(\mathbf{k}-\mathbf{q})-\delta_{\ell_{v}\ell}\phi_{\nu\sigma}(\mathbf{k})$. The calculation of the direct interaction $U_{\nu\tilde{\nu}}^{\nu^{\prime}\tilde{\nu}^{\prime}}$ is straightforward while evaluating the exchange interaction $V_{\nu\tilde{\nu}}^{\nu^{\prime}\tilde{\nu}^{\prime}}$ is rather involved and is detailed in SM [53]. We remark that, the result in Eq.(9) is for intravalley interaction and only valid for the low-density regime.
@@
Figure 2: (a) Exciton energy $\mathcal{E}$ ($\sigma=\uparrow$) and twisting angle $\theta$ dependence. The points are color coded by its light-matter coupling strength $g_{c}$. The hybridized excitons weakly interact with light are shown in gray dots. The red and blue dashed lines are the A and B exciton energy with $\theta=60^{\circ}$. The vertical dashed line is given by $\frac{\Delta\mathbf{K}^{2}}{2m_{v}}=t_{v}$ where the energy separation between folded bands is comparable to interlayer tunneling energy (Inset) the moiré exciton spectrum is calculated by considering the Umklapp scattering that hybridizes the exciton in zone 1 (purple area) and zone 2 (gray area). In the weak Umklapp region, the interlayer and intralayer exciton hybridized with each mBZ separately. (b) $\theta$-dependence of the exciton-exciton interaction. (Inset) the interaction (wavy black curves) between interlayer excitons with electron exchange (red dashed curves).
@@
$\mathcal{H}_{\mathrm{xp}}\!=\!\omega_{c}c^{\dagger}_{0}c_{0}+\sum_{\bar{\alpha}}(\mathcal{E}^{\bar{\alpha}}_{\sigma}\!+\tfrac{1}{L^{2}}g^{\bar{\alpha}}_{X}\hat{x}^{\dagger}_{\bar{\alpha}\sigma}\hat{x}_{\bar{\alpha}\sigma})\hat{x}^{\dagger}_{\bar{\alpha}\sigma}\hat{x}_{\bar{\alpha}\sigma}+\mathcal{H}_{\mathrm{sc}}$ (10)
where the X-X interaction is $g^{\bar{\alpha}}_{X}\approx L^{2}\Delta^{\bar{\alpha}\bar{\alpha}}$. The exciton-photon interaction is $g_{c}^{\bar{\alpha}}=\frac{1}{2}\sum_{\alpha\ell}\Omega^{\ell}_{\alpha}\bar{C}^{\bar{\alpha},\ell\ell}_{\alpha\sigma}$, where $\Omega^{\ell}_{\alpha}=2\sum_{\mathbf{k}}d^{\ell}_{cv}\phi^{\ell\ell}_{\alpha}(\mathbf{k})$ is the Rabi splitting of the exciton in $\ell$-th layer and state $\alpha$.
@@
[…]
@@
In our analysis, we focus on the nonlinear response and X-X interaction strength shown in Fig. 2b. In the plot we keep only the dominant contributions from Eq. (9) with $\bm{g}^{\prime}=\bm{g}$ for simplicity, since the Coulomb scattering processes with nonzero momentum transfer between the excitons are rapidly suppressed as $\theta\gtrsim 0.5^{\circ}$. First of all, we find an enhancement in the nonlinear interaction at the small-$\theta$ regime indicating the strong influence of the band hybridization due to the Umklapp processes. We find strong attractive nonlinearity for hIX while the nonlinear interaction for hX remains repulsive. Intriguingly, these results correlate with measurements in MoS2 bilayers, where redshifts were observed [30]. The emergence of attractive nonlinearity may be understood as a manifestation of strong anisotropic nature of 2D bilayers. This leads to a weaker interlayer e-h interaction as compared to the intralayer e-e and h-h interaction (Fig. 2b, inset). Hence, the total nonlinearity from each of these interacting channels does not cancel leading to stronger attractive nonlinearity. This is different from the monolayer exciton where large cancellations [63, 64, 65] between these channels result in a weaker repulsive nonlinearity ($\sim 0.5$ nm2eV). This anisotropic electronic property is unique for 2D materials that enable the realization of attractive interaction between dipolar excitons.
We also note that the strong attractive X-X interaction between hIX in the small $\theta$ regime comes from the exchange scattering involving the excitonic modes in Zone 2 (Fig. 2a, inset). For instance, in the scattering processes with $\bm{g}\neq\tilde{\bm{g}}$ (Fig. 1d), the second term in the last line in Eq.(9) (all particle exchange process, repulsive) is strongly suppressed due to large momentum transfer ($\mathbf{q}=\bm{g}-\tilde{\bm{g}}$), while the exchange interaction $V_{\nu\tilde{\nu}}^{\nu^{\prime}\tilde{\nu}^{\prime}}$ and $V_{\nu\tilde{\nu}}^{\tilde{\nu}^{\prime}\nu^{\prime}}$ remain large. This makes a striking difference between moiré bilayer and untwisted bilayer ($\theta=0$ and $60^{\circ}$) where the scattering processes only take place within Zone 1 with $\bm{g}=\tilde{\bm{g}}=0$. Furthermore, in our calculation, we find weak repulsive direct interaction for hIX in a homobilayer with $\theta\approx 0$ (see SM[53]). This may be expected since a large electrical dipole moment is forbidden due to the approximate inversion symmetry. As a result, the dominant contribution to $g_{X}$ in Fig. 2b comes from the (attractive) exchange interactions. However, this can change for heterobilayers or electrically-biased samples.
@@
In the presence of electric field $\mathbf{E}_{z}$, the dipole moment of the interlayer exciton couples to $\mathbf{E}_{z}$ giving rise to the energy Stark shift. This changes the hybridization content of hX and hIX leading to electrically-tunable optical properties. The Stark effect can be incorporated into Eq. (7) by modifying the interlayer exciton energy $E_{\nu\sigma}(\bar{\mathbf{Q}})\to E_{\nu\sigma}(\bar{\mathbf{Q}})+\mathbf{E}_{z}\cdot\bm{\mu}^{s}$, where the dipole moment $\bm{\mu}^{s}=\pm\mu_{z}\hat{\mathbf{z}}$ for interlayer exciton with $\mu_{z}\sim 5$eÅ [66, 29], and $\mu_{z}=0$ for intralayer exciton. In Figs. 3a and 3b, we calculate the absorption $\Gamma(\mathcal{E})$. In $\Gamma(\mathcal{E})$, we also take into account the nonlinear energy shift due to the background excitons with density $n_{\bar{\alpha}}=\langle\hat{x}^{\dagger}_{\bar{\alpha}\sigma}\hat{x}_{\bar{\alpha}\sigma}\rangle/L^{2}$ (pump-power dependent)[53]. In Figs. 3a, two brightest branches independent of $E_{z}$ correspond to the hybridized A and B intralayer excitons. The other dimmer branches with strong response to $E_{z}$ are the hIXs. Following one hIX branch indicated by the red and blue dots in Figs. 3a and 3b, we can see that the absorption peaks are enhanced as the hIX approaches the hX branches since hIX gains more intralayer exciton component. Furthermore, we also observe that the hIX nonlinear redshift turns into a blueshift as $E_{z}$ changes. This is partly due to the repulsive nonlinearity from the intralayer exciton component, and another contribution is coming from the polarization of hIX in the high-field regime. These results in Figs. 2 and 3 demonstrate that the nonlinear optical response of moiré material is electrically tunable.
@@
Next, we plot the polaritonic spectrum in Fig. 3(c) by evaluating the photonic density of states $\rho_{c}(\omega)$ of Eq. (10), which is proportional to the cavity transmission [53]. Tuning the hIX energy closer to the hX energy at $E_{z}=0.5$ eV/nm, we enhance the light-matter coupling of hIX leading to the larger Rabi splitting. This also changes the hIX attractive nonlinearity (red arrows) at $E_{z}=0$ to a repulsive nonlinearity (blue arrows) with blueshift at $E_{z}=0.5$ V/nm.
@@
Conclusion.—We developed a microscopic theory for hybridized moiré excitons in twisted bilayers. We revealed that hybridization between layers can enhance significantly the nonlinearity of moiré excitons and polaritons, stemming from Umklapp processes for small twisting angles. Intriguingly, we find attractive nonlinear interaction for hIX, which can be tuned into repulsive by applying an external electric field. This makes moiré polariton lattices at small twisting angles an excellent platform for studying many-body effects.
@@
Note that in this study we limited ourselves to not-too-large $\theta$ regime, since the hybridization between folded bands in the mini Brillouin zone is weak. On the other hand, we do not investigate $\theta\lesssim 0.5^{\circ}$, as this is typically prevented by the lattice reconstruction [67, 68, 69, 70, 71] and also requires accounting for more Umklapp scatterings beyond those pictured in Fig. 2a. We also stress that the nonlinear properties in our study are valid for low density regimes, and studying higher-order correction is an interesting avenue for future research. This will allow describing a crossover to the strongly correlated regime close to the metal-insulator transition [72, 73, 74, 75].